# TRIM: Token-wise Attention-Derived Saliency for Data-Efficient Instruction Tuning

### v1

@@ 3.2 Stage II: Candidate Scoring and Selection
If a token class $t_{j}$ does not appear in $\mathcal{T}_{\text{val}}$, it will not have a corresponding fingerprint in $\mathcal{F}$. For these unseen classes, we find the nearest fingerprinted class in the input-embedding space via cosine similarity and apply a penalty factor $\lambda\in(0,1]$ to down-weight its contribution. Let $E$ denote the input-embedding matrix of $\mathbf{M}_{\text{warmup}}$ and $e_{t}$ be the unit-normalized embedding for class $t$. The nearest fingerprinted class is:
@@
The final score for a candidate $c$ is aggregated over its token set $M(c)$ using a robust pooling function that combines the mean and maximum of its token scores:
$S(c)=w_{\mu}\tfrac{1}{|M(c)|}\!\sum_{j\in M(c)}\!s_{j}+w_{m}\max_{j\in M(c)}s_{j}$ (13)
@@
Finally, we rank all candidate examples $c\in\mathcal{S}$ by their scores $S(c)$ and select the top-ranking examples to form the coreset $\mathcal{C}$. Since this scoring stage only requires a single forward pass per candidate, its runtime is linear in the corpus size once the fingerprints are constructed. For detailed pseudocode, please refer to Appendix˜B.
@@ 4 Experiments
We assess TRIM on three axes: (i) budgeted accuracy across general-domain multiple-choice reasoning benchmarks (Section˜4.1) and a low-overlap out-of-domain math benchmark (Section˜4.2); (ii) cross-model/scale transfer (Section˜4.3); and (iii) robustness to length bias (Section˜4.4).
@@
We select from a large candidate pool of 270k instruction-tuning examples (Dolly (Conover et al., 2023), CoT (Wei et al., 2022), Oasst1 (Köpf et al., 2023), Flan_v2 (Longpre et al., 2023)) to fine-tune models for general-domain multiple-choice reasoning benchmarks (CommonsenseQA (Talmor et al., 2019), SocialIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019)) and a mathematical reasoning benchmark (GSM8K (Cobbe et al., 2021)). We use Llama-3.2-1B for the main comparisons and Llama-2-7B (Touvron et al., 2023) for GSM8K. For transferability, we evaluate how coresets selected by a Llama-3.2-1B scorer are applied to larger or different architectures, specifically Llama-3.1-8B (Grattafiori et al., 2024) and Mistral-7B (v0.3) (Jiang et al., 2023). All models are trained with LoRA (Hu et al., 2022), and selection of candidate samples is guided by 5–10 target validation samples per task. The full setup and hyperparameter details are provided in Appendices DE, and F.
@@
We group coreset methods by mechanism and defer the details to Appendix C. (i) Heuristics: Random, lexical retrieval (BM25 (Robertson et al., 2009)), and $N$-gram importance resampling (DSIR (Xie et al., 2023)). (ii) Forward-only training dynamics: CLD (correlation of loss differences) (Nagaraj et al., 2025) and S2L (loss trajectory clustering) (Yang et al., 2024). (iii) Representation-based: RDS (Hanawa et al., 2021). (iv) Gradient-based: LESS (Xia et al., 2024) and TAGCOS (Zhang et al., 2025b). All methods share a standardized warmup-and-selection pipeline (Appendix F). We warm up Llama-3.2-1B on a random 5% subset of the source dataset using LoRA. TRIM/TAGCOS score candidates using the final warmup checkpoint, LESS/RDS aggregate across checkpoints, and CLD/S2L use trajectories collected during full-dataset training. For context, we also report Pretrained and Full-data baselines.
@@ 4.1 Performance Comparison
We evaluate all methods at a fixed 5% budget. Each method selects its coreset from the candidate pool, which is then used to fine-tune a fresh Llama-3.2-1B model with LoRA, initialized from the pretrained checkpoint (i.e., not the warmed-up scorer).
Coreset Method CommonsenseQA SocialIQA HellaSwag Average Mean
@@
Random 34.05 (1.2) 43.84 (0.1) 48.21 (0.1) 42.03
DSIR (Xie et al., 2023) 37.16 (0.7) 44.38 (0.5) 48.75 (0.2) 43.43
S2L (Yang et al., 2024) 34.10 (0.1) 43.32 (0.2) 48.57 (0.2) 41.99
RDS (Hanawa et al., 2021) 36.16 (1.1) 43.77 (0.1) 48.43 (0.3) 42.79
CLD (Nagaraj et al., 2025) 33.10 (0.4) 43.25 (0.5) 48.35 (0.1) 41.56
LESS (Xia et al., 2024) 39.10 (0.7) 44.52 (0.3) 49.01 (0.1) 44.21
TAGCOS (Zhang et al., 2025b) 34.72 (0.8) 43.70 (0.2) 48.76 (0.1) 42.39
@@
Results and Takeaways.
As summarized in Table˜1, TRIM attains the highest mean accuracy of 45.37%, exceeding the next-best method, LESS (44.21%), by over one point. Notably, TRIM’s 5% coresets are the only ones that surpass full-data fine-tuning on two tasks, SocialIQA (46.26% vs. 45.04%) and HellaSwag (49.08% vs. 48.55%).
Figure 2: TRIM vs. top baselines on Llama-3.2-1B across coreset budgets for CommonsenseQA (left) and SocialIQA (right). TRIM maintains a consistent advantage and, on SocialIQA, exceeds the full-data baseline with coresets as small as 1%.
Beyond the 5% budget, fig.˜2 shows TRIM consistently outperforming LESS and TAGCOS across coreset sizes. On SocialIQA, it surpasses full-data at 1% and peaks at 2.5%; on CommonsenseQA, it exceeds full-data at larger budgets ($20\%$).
While heuristic lexical methods (BM25, DSIR) track surface overlap; training-dynamics (CLD, S2L), representation-based (RDS), and gradient-based (LESS, TAGCOS) remain sample-level and thus coarse. On the other hand, TRIM’s token-centric scoring targets critical tokens, enabling a more precise selection. We provide an ablation study over our hyperparameters in Table G.
4.2 Low-Overlap Out-Of-Domain: Mathematical Reasoning
We stress test TRIM’s core hypothesis: token-level representational matching can surface training data that imparts structural reasoning skills even without topical overlap. Concretely, we ask each method to select a coreset for GSM8K from our general-purpose candidate instruction-tuning pool (not curated for math). We then fine-tune Llama-2-7B on the selected data and evaluate on GSM8K. This setting is challenging; fine-tuning on the entire corpus yields only 30.25%, so success requires identifying examples with latent structures (e.g., stepwise logic, numerical manipulation) rather than superficial topical cues. We include three strong sample-wise baselines: LESS, S2L, and TAGCOS.
@@ 4.2 Low-Overlap Out-Of-Domain: Mathematical Reasoning
LESS (Xia et al., 2024) 17.34 (1.2) 20.72 (0.4)
S2L (Yang et al., 2024) 16.48 (0.4) 18.55 (0.3)
TAGCOS (Zhang et al., 2025b) 17.23 (0.2) 18.72 (0.4)
@@
Table 2: Out-of-domain data selection for GSM8k on Llama-2-7B. The table shows exact-match accuracy (%) for coresets of 1% and 5% selected from a non-math corpus. All scores are the mean of 3 seeds.
Results and Takeaways.
As shown in Table˜2, a 5% TRIM coreset reaches 29.52%, nearly matching full-data fine-tuning (30.25%) and outperforming the next best method (LESS) by $\sim$8.8%. Even at a 1% budget, TRIM achieves 22.33% accuracy, remaining well ahead of alternatives. This stems from strong structural fidelity: by matching token-level fingerprints, rather than sample-level losses or gradients, TRIM retrieves examples that convey a style of reasoning (e.g., enumerations, logical connectives, stepwise instructions). These signals act as transferable proxies for chain-of-thought structure in GSM8K, yielding a higher performance that sample-level metrics fail to capture.
@@ 4.3 Coreset Transferability
We assess whether a small scorer can curate data that transfers to larger models. Using Llama-3.2-1B as the scorer, we select a single 5% coreset and reuse it to fine-tune two larger target models: Llama-3.1-8B (same model family) and Mistral-7B (v0.3) (different model family). We compare (i) TRIM-Transfer, the coreset chosen by the 1B scorer, and (ii) TRIM-Oracle, a coreset selected using the target model itself as the scorer. The oracle serves as an in-model reference point.
Llama-3.1-8B Mistral-7B (v0.3)
Coreset Method CSQA SIQA HS Avg. CSQA SIQA HS Avg.
@@
Results and Takeaways.
As shown in Table˜3, coresets selected by the 1B scorer transfer with high fidelity: on Llama-3.1-8B the transferred coreset reaches 64.17% vs. the oracle’s 63.48%, and on Mistral-7B it again edges the oracle (65.39% vs. 64.83%). Thus, a single coreset selected once by a small model can match, or even surpass, a per-target oracle. This strong transfer suggests that TRIM’s token-level fingerprints capture model-agnostic structure that generalizes across scale and architecture, enabling a lightweight scorer to curate high-quality coresets for large targets without sacrificing accuracy.
@@ 4.4 Mitigating Length Bias
A common failure mode in coreset selection is a bias toward shorter sequences. Sample-level heuristics can be confounded by sequence length, e.g., influence scores derived from gradient magnitudes tend to favor short examples with concentrated signals over longer prompts that contain richer reasoning (Xia et al., 2024). We test whether TRIM exhibits this bias by analyzing the length distribution of coresets selected for GSM8K by TRIM, LESS, and TAGCOS.
Figure 3: Length distribution of selected coresets for GSM8K. The histogram shows the percentage of selected samples by length bucket. Sample-level methods (LESS, TAGCOS) skew short, whereas TRIM selects a broader distribution with a higher mean length.
Results and analysis.
As shown in Figure˜3, LESS and TAGCOS skew toward shorter samples (mean lengths 323 and 259 tokens). In contrast, TRIM selects a broader distribution with many longer, more complex examples; its mean length is 446 tokens, $\sim$38% longer than LESS and $\sim$72% longer than TAGCOS. Since sample-level scoring can conflate brevity with quality, TRIM’s token-centric approach is effectively length-agnostic: by elevating important tokens wherever they occur, it yields diverse, informative coresets.
4.5 Task-Adaptive Data Selection
We analyze the distribution of selected examples across GSM8K and CommonsenseQA, as shown in Figure 4. TAGCOS maintains nearly identical selection patterns for both tasks, demonstrating its inability to adapt to task-specific requirements. In contrast, TRIM exhibits strong task-adaptive behavior: for CommonsenseQA, it predominantly selects from Flan_v2, leveraging its diverse instruction-following examples well-suited for question answering, while for GSM8K, it shifts toward the CoT dataset, prioritizing step-by-step reasoning essential for mathematical problem solving. Notably, LESS struggles to identify the importance of CoT data for GSM8K, while still favoring Flan_v2. This demonstrates that effective data selection requires not just quality assessment, but also the ability to recognize task-specific data characteristics, a capability that TRIM successfully captures. For more results, refer to Appendix˜H.
Figure 4: Distribution of selected training examples across data subsets for GSM8K and CommonsenseQA.
@@ 5 Computational Cost of Data Selection
We analyze the asymptotic cost of the selection stage (model preparation + candidate scoring). Let $N$ be the number of candidate examples and $T$ the number of training epochs used for model preparation (“warmup”). Let $f$ denote the FLOPs of a single forward pass of the scoring model. Following convention, a backward pass costs $\approx 2f$, so a full training step (forward + backward) is $\approx 3f$. We denote by $\gamma\in(0,1]$ the fraction of the dataset used for warmup and by $C$ the number of checkpoints whose states/representations are used during scoring. For a fair comparison, we assume all methods use a scoring model with forward-pass cost $f$. The cost to process the (small) target validation set of size $Q$ is negligible since $N\gg Q$ (in our setup, N>270N\!>\!270k and $Q\in\{5,10\}$).
Coreset Method Total selection cost
S2L (Yang et al., 2024) $\mathcal{O}(3fNT\;+\;fNC)$
LESS (Xia et al., 2024) $\mathcal{O}(3f\,\gamma NT\;+\;3fNC)$
TAGCOS (Zhang et al., 2025b) $\mathcal{O}(3f\,\gamma NT\;+\;3fN)$
RDS (Hanawa et al., 2021) $\mathcal{O}(3f\,\gamma NT\;+\;fNC)$
@@
Table 4: Asymptotic selection cost under a common scoring model (forward cost $f$). The first term is model preparation (warmup), the second is candidate scoring.
With a fixed scoring model size, the drivers are (i) whether backward passes are required during scoring and (ii) how many checkpoints $C$ are used. LESS is most expensive at scoring time: it uses gradient features at multiple checkpoints, incurring a backward pass for each candidate at each checkpoint, yielding a $\mathcal{O}(3fNC)$ term in addition to warmup $\mathcal{O}(3f\,\gamma NT)$. Forward-only methods that do not compute gradients at scoring time (e.g., S2L, RDS) avoid the $3f$ factor at scoring but still scale linearly in $C$ via $\mathcal{O}(fNC)$. Importantly, S2L (Yang et al., 2024) must train on all $N$ candidates to record per-sample loss trajectories (the signal it clusters), so its preparation term is $\mathcal{O}(3fNT)$ rather than $\mathcal{O}(3f\,\gamma NT)$. TAGCOS scores at a single checkpoint (C=1C{=}1) but requires per-candidate gradients once, giving $\mathcal{O}(3fN)$ after warmup. In contrast, TRIM scores each candidate with a single forward pass at one checkpoint (C=1C{=}1), producing $\mathcal{O}(fN)$ after warmup. Consequently, TRIM is asymptotically the most efficient among the compared methods for large $N$, and its benefits increase when baselines require multiple checkpoints ($C\gg 1$) or gradient computation during scoring.

### v2

@@ 3.2 Stage II: Candidate Scoring and Selection
If a token class $t_{j}$ does not appear in $\mathcal{T}_{\text{val}}$, it will not have a corresponding fingerprint in $\mathcal{F}$. For such tokens, we map $t_{j}$ to the closest fingerprinted class in the model’s input-embedding space (by cosine similarity), and down-weight its contribution by a penalty factor $\lambda\in(0,1]$. We precompute and cache this nearest-class mapping once per tokenizer/model (Appendix C). Let $E$ denote the input-embedding matrix of $\mathbf{M}_{\text{warmup}}$ and let $e_{t}$ be the unit-normalized embedding for class $t$. The fallback mapping is:
@@
The final score for a candidate $c$ is aggregated over its token set $\mathcal{J}(c)$ using a robust pooling function that combines the mean and maximum of its token scores:
$S(c)=w_{\mu}\tfrac{1}{|\mathcal{J}(c)|}\!\sum_{j\in\mathcal{J}(c)}\!s_{j}+w_{m}\max_{j\in\mathcal{J}(c)}s_{j}$ (13)
@@
Finally, we rank all candidate examples $c\in\mathcal{S}$ by their scores $S(c)$ and select the top-ranking examples to form the coreset $\mathcal{C}$. Since this scoring stage only requires a single forward pass per candidate, the runtime is linear in the corpus size once the fingerprints are constructed. For detailed pseudocode, please refer to Appendix B.
@@ 4 Experiments
We evaluate TRIM along four axes: (i) budgeted accuracy on general-domain reasoning benchmarks and a challenging downstream math task (Sections 4.1 and 4.2); (ii) cross-model/scale transfer (Section 4.3); (iii) robustness to length bias (Section 4.4); and (iv) an in-domain setting to study coreset selection when there are no external target examples, with the subset selection relying on the candidate pool $\mathcal{S}$ itself (Section 4.5). Additional hyperparameter ablations, experiments on sensitivity to target samples, as well as performance on additional models and benchmarks, are provided in the Appendices G, H, I, J, and N.
@@
For all experiments, we follow the warmup and selection protocol of Xia et al. (2024). We first fine-tune a scorer on a random 5% subset of the candidate pool $\mathcal{S}$ using LoRA (Hu et al., 2022), saving intermediate checkpoints. We then score all candidates in $\mathcal{S}$ using this warmed-up scorer model and select a budgeted coreset. Finally, we fine-tune a fresh target model initialized from the pretrained checkpoint on the selected coreset and report downstream performance.
General instruction-pool setting.
Unless stated otherwise, we select from a large candidate pool of 270k instruction-tuning examples (Dolly (Conover et al., 2023), Oasst1 (Köpf et al., 2023), Flan_v2 (Longpre et al., 2023), and CoT (Wei et al., 2022)). Our target tasks are MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2023), TyDiQA (Clark et al., 2020), and GSM8K (Cobbe et al., 2021). To unify heterogeneous sources, we format all datasets into a chat-style schema following the instruction-tuning pipeline of Wang et al. (2023). We evaluate with lm-evaluation-harness (Sutawika et al., 2023) using fixed few-shot prompting: 5-shot MMLU, 3-shot BBH, 1-shot TyDiQA, and 8-shot GSM8K with CoT. Additional dataset and evaluation details are in Appendices E and F.
Models.
We use Llama-2-7B (Touvron et al., 2023) as the base model for all main experiments. To study transferability, we test whether coresets selected using a Llama-2-7B scorer transfer to Llama-2-13B (Touvron et al., 2023) and Mistral-7B (v0.3) (Jiang et al., 2023). Data selection is guided by a small target validation set per benchmark, where the number of validation examples is determined by the benchmark’s few-shot configuration following Xia et al. (2024). The training details are provided in the Appendix G.
@@
We group coreset methods by mechanism and defer details to Appendix D. (i) Heuristics: Random, lexical retrieval (BM25 (Robertson et al., 2009)), and $N$-gram importance resampling (DSIR (Xie et al., 2023)). (ii) Forward-only training dynamics: CLD (correlation of loss differences) (Nagaraj et al., 2025) and S2L (loss-trajectory clustering) (Yang et al., 2024). (iii) Representation-based: RDS (Hanawa et al., 2021). (iv) Gradient-based: LESS (Xia et al., 2024) and TAGCOS (Zhang et al., 2025b). TRIM/TAGCOS score candidates using the final warm-start checkpoint, LESS/RDS aggregate scores across checkpoints, and CLD/S2L rely on trajectories collected during full-dataset training.
@@ 4.1 Performance Comparison
Table 1 compares the performance of various data selection methods at a fixed 5% budget. TRIM achieves the best overall performance, with a mean accuracy of 48.56%, showing comparative performance to the strongest baseline, LESS (48.27%), while requiring substantially less computation. Notably, TRIM and LESS are the only methods whose 5% coresets surpass full-data fine-tuning on TyDiQA.
Table 1: Accuracy (%) for coreset methods on Llama-2-7B with a 5% coreset. Results are means over 3 seeds (standard deviations in parentheses). The final column is the macro-average across tasks.
Coreset Method MMLU BBH TyDiQA Average Mean
@@
Random 45.84 (0.4) 37.13 (0.4) 52.45 (0.6) 45.14
BM25 (Robertson et al., 2009) 46.12 (0.0) 37.4 (0.0) 52.70 (0.0) 45.41
DSIR (Xie et al., 2023) 45.73 (0.1) 34.70 (0.3) 45.62 (0.5) 42.02
S2L (Yang et al., 2024) 46.70 (0.2) 37.24 (0.4) 52.74 (0.3) 45.56
RDS (Hanawa et al., 2021) 45.27 (0.6) 35.13 (0.5) 46.10 (0.2) 42.17
CLD (Nagaraj et al., 2025) 46.13 (0.2) 36.12 (0.3) 45.58 (0.5) 42.61
LESS (Xia et al., 2024) 49.23 (0.3) 39.45 (0.2) 56.13 (0.4) 48.27
TAGCOS (Zhang et al., 2025b) 48.12 (0.2) 38.24 (0.3) 53.45 (0.4) 46.60
@@
Figure 2: TRIM vs. top baselines on Llama-2-7B across coreset budgets for MMLU, BBH, and TyDiQA. TRIM maintains a consistent advantage and, on TyDiQA, exceeds the full-data baseline with coresets as small as 1%.
Beyond the 5% budget, Figure 2 shows TRIM consistently matching or improving upon strong baselines (LESS, TAGCOS) across coreset sizes. On MMLU and TyDiQA, TRIM remains competitive across all budgets and exceeds the full-data reference once moderate budgets are available; on BBH, it maintains a stable advantage over TAGCOS and tracks LESS closely as the budget increases. Overall, heuristic lexical methods (BM25, DSIR) primarily track surface overlap, while training-dynamics (CLD, S2L), representation-based (RDS), and gradient-based (LESS, TAGCOS) selectors operate at the sample level and are therefore coarser. In contrast, TRIM’s token-centric scoring emphasizes task-relevant tokens, enabling a precise selection.
4.2 Challenging Downstream Tasks
Table 2: Challenging downstream data selection for GSM8K on Llama-2-7B. The table shows exact-match 8-shot accuracy (%) for coresets of 1% and 5% selected from a non-math corpus. Results are means over 3 seeds (standard deviations in parentheses).
@@ 4.2 Challenging Downstream Tasks
S2L 16.48 (0.4) 18.55 (0.3)
LESS 17.34 (1.2) 20.72 (0.4)
TAGCOS 17.23 (0.2) 18.72 (0.4)
@@
We rigorously evaluate the role of the token-centric approach in challenging downstream-task settings, where the target task has a low topical overlap with the candidate pool. Concretely, we ask each method to select a coreset for GSM8K from the same general-purpose instruction-tuning pool (not curated for math), then fine-tune on the selected data and evaluate on GSM8K. This setting is challenging: fine-tuning on the full pool yields only 30.25%, so success requires identifying examples with latent structure (e.g., stepwise logic, numerical manipulation) rather than superficial topical cues. As shown in Table 2, a 5% TRIM coreset reaches 29.52%, nearly matching full-data fine-tuning and outperforming the next best method (LESS) by $\sim$8.8%. Even at a 1% budget, TRIM achieves 22.33% accuracy, remaining well ahead of alternatives. This stems from strong structural fidelity: by matching token-level fingerprints rather than sample-level losses or gradients, TRIM retrieves examples that convey a style of reasoning (e.g., enumerations, logical connectives, stepwise instructions). These signals act as transferable proxies for chain-of-thought structure in GSM8K, yielding gains that sample-level metrics fail to capture.
@@ 4.3 Coreset Transferability
Table 3: Cross-Model and Cross-Scale Transferability of Coresets Selected by a 7B Scorer. Results are means over 3 seeds (standard deviations in parentheses). Bold numbers denote the best performing selected subset. Underlined numbers denote that the selected subset outperformed the full-data fine-tuning.
Llama-2-13B Mistral-7B (v0.3)
Coreset Method MMLU BBH TyDiQA MMLU BBH TyDiQA
@@
TRIM-Transfer (from 7B) 54.50(0.2) 49.83(0.6) 56.70(0.3) 60.40(0.8) 55.50(0.2) 60.43(0.4)
TRIM-Oracle (in-model) 54.10 (0.6) 50.02(0.7) 54.12(0.2) 61.72(0.3) 56.12(0.4) 59.83(0.2)
We assess whether a scorer model can curate a coreset that transfers across scale and architecture. Using Llama-2-7B as the scorer, we select a single 5% TRIM coreset and reuse it to fine-tune two target models: Llama-2-13B (larger model within the same family) and Mistral-7B (v0.3) (a different architecture). We compare (i) TRIM-Transfer, where selection is performed once by the 7B scorer, against (ii) TRIM-Oracle, where the target model selects its own coreset (an in-model reference) in Table 3. On Llama-2-13B, TRIM-Transfer matches or exceeds the in-model oracle on MMLU and substantially improves TyDiQA, while remaining competitive on BBH (49.83 vs. 50.02). On Mistral-7B, TRIM-Transfer remains robust across all three benchmarks, outperforming full-data fine-tuning and approaching the oracle (TyDiQA: 60.43 vs. 59.83), though the oracle is strongest on MMLU and BBH. Overall, these results suggest that TRIM’s token-level fingerprints capture transferable, model-agnostic structure, enabling a single scorer to select high-quality coresets that generalize across scale and architecture.
@@ 4.4 Mitigating Length Bias
A common failure mode in coreset selection is a bias toward shorter sequences. Sample-level scores can be confounded by length; for example, influence estimates derived from gradient magnitudes may favor short examples with concentrated signals over longer prompts containing richer reasoning (Xia et al., 2024). We analyze this effect by comparing the length distributions of coresets selected for GSM8K by TRIM, LESS, and TAGCOS. As shown in Figure 3, LESS and TAGCOS skew short (mean lengths 323 and 259 tokens). In contrast, TRIM selects a broader distribution with more long, complex examples (mean length 446 tokens), $\sim$38% longer than LESS and $\sim$72% longer than TAGCOS. This behavior is consistent with TRIM’s token-centric scoring: relevance is assigned to salient tokens, and pooling aggregates token matches without directly rewarding shorter sequences.
Figure 3: Length distribution of selected coresets for GSM8K. The histogram shows the percentage of selected samples by length bucket. Sample-level methods (LESS, TAGCOS) skew short, whereas TRIM selects a distribution with a higher mean length.
4.5 In-Domain Data Selection Without Target Examples
Table 4: Exact-match 0-shot accuracy (%) on GSM8K and SVAMP for coreset methods on Llama-2-7B fine-tuned on a $50\%$ coreset selected from MathInstruct. TRIM achieves accuracy comparable to prior in-domain selectors while using substantially lower selection compute (see Section 5). Results are means over 3 seeds (standard deviations in parentheses).
Coreset Method GSM8K SVAMP
[…]
Random 48.16 (0.5) 63.30 (1.2)
S2L 52.45 (0.1) 65.30 (0.2)
LESS 52.10 (0.1) 65.45 (0.2)
TAGCOS 50.20 (0.2) 64.15 (0.4)
[…]
So far, our selection setups assume access to a separate target/validation set (from the downstream task) to instantiate $\mathcal{T}_{\text{val}}$ for validation and fingerprinting. Here, we consider a stricter in-domain regime where such target examples are unavailable: selection must use only the candidate pool, aiming to preserve the pool distribution.
We instantiate this setting with the math-focused MathInstruct corpus (Yue et al., 2024) as the candidate pool $\mathcal{S}$, formatted in the Alpaca chat schema (Chen et al., 2024b), and construct an in-domain reference set $\mathcal{T}_{\text{val}}$ by sampling m=20m{=}20 examples uniformly from each of the 13 MathInstruct sources ($|\mathcal{T}_{\text{val}}|{=}260$). Unlike prior sections where $\mathcal{T}_{\text{val}}$ is disjoint from $\mathcal{S}$, here $\mathcal{T}_{\text{val}}\subset\mathcal{S}$ by construction. We then follow the standard protocol, a warmed-up Llama-2-7B (trained on 5% random MathInstruct) scores candidates in $\mathcal{S}$; we select a 50% coreset; and fine-tune a fresh pretrained Llama-2-7B on the selected subset. Table 4 reports 0-shot exact-match on GSM8K and SVAMP under the MAmmoTH pipeline. TRIM matches prior works such as S2L/LESS [^1] To ensure fairness, $\mathcal{T}_{\text{val}}$ is the same for both TRIM and LESS. within $1\%$ while being substantially more compute-efficient (see Section 5). Additional details, ablations, and extended results are in Appendix M.
@@ 5 Computational Cost of Data Selection
We analyze the cost of the selection stage (model preparation + candidate scoring) and complement the asymptotic analysis with measured wall-clock runtimes. Let $N$ be the number of candidate examples and $T$ the number of training epochs used for model preparation (i.e., warmup). Let $f$ denote the FLOPs of a single forward pass of the scoring model. Following convention, a backward pass costs $\approx 2f$, so a training step (forward + backward) costs $\approx 3f$. We write $\gamma\in(0,1]$ for the fraction of the candidate pool used during warmup and $C$ for the number of checkpoints whose states/representations are used during scoring. For a fair comparison, we assume all methods use a scoring model with forward-pass cost $f$. Finally, processing the target validation set is negligible compared to scoring the candidate pool since $N\gg Q$ (in our setup, N>270N{>}270k while $Q\leq 285$).
With a fixed scoring model size, the main drivers are (i) whether scoring requires backward passes and (ii) how many checkpoints $C$ are used. LESS is most expensive at scoring time: it aggregates gradient features across multiple checkpoints, incurring a backward pass per candidate per checkpoint and yielding a $\mathcal{O}(3fNC)$ scoring term in addition to warmup $\mathcal{O}(3f\,\gamma NT)$. Forward-only methods that avoid gradients at scoring time (e.g., S2L, RDS) remove the $3f$ factor during scoring but still scale linearly with $C$ via $\mathcal{O}(fNC)$. Notably, S2L must train on all $N$ candidates to record per-sample loss trajectories (the signal it clusters), so its preparation term is $\mathcal{O}(3fNT)$ rather than $\mathcal{O}(3f\,\gamma NT)$. TAGCOS scores at a single checkpoint (C=1C{=}1) but requires per-candidate gradients once, giving $\mathcal{O}(3fN)$ after warmup. In contrast, TRIM scores each candidate with a single forward pass at one checkpoint (C=1C{=}1), producing $\mathcal{O}(fN)$ after warmup. Consequently, TRIM is asymptotically the most efficient among the compared methods for large $N$, and its benefits grow when baselines require multiple checkpoints ($C\gg 1$) or gradient computation during scoring.
To ground the above analysis, we measure end-to-end selection time (warmup + scoring) under a shared setup: a single NVIDIA H200 GPU, a common candidate pool of 270k instruction-tuning examples (from Dolly, CoT, Oasst1, Flan_v2), and a Llama-2-7B scorer selecting a 5% coreset for MMLU. All methods use the same warmup recipe, yielding a warmup time of 1.41 hours, except S2L, which requires full-data training to collect loss trajectories and therefore incurs a substantially larger preparation cost (27.96 hours). Table 5 reports the asymptotic selection cost alongside the measured total runtime. Overall, the measured runtimes mirror the asymptotic analysis: TRIM scores candidates using only forward passes (and a single checkpoint), making it considerably faster than methods that rely on gradients during scoring (LESS, TAGCOS) or full-data trajectory collection (S2L).
Table 5: Selection cost and measured runtime (5% coreset). Theoretical cost is for the selection stage (warmup + scoring) under a common scoring model with forward-pass cost $f$. The runtime column reports the total wall-clock time (hours).
Coreset
Method
Asymptotic
Selection Cost
Total Runtime
(in Hours)
S2L $\mathcal{O}(3fNT\;+\;fNC)$ 87.95
RDS $\mathcal{O}(3f\,\gamma NT\;+\;fNC)$ 20.48
LESS $\mathcal{O}(3f\,\gamma NT\;+\;3fNC)$ 35.76
TAGCOS $\mathcal{O}(3f\,\gamma NT\;+\;3fN)$ 14.16